OpenStudy (anonymous):

help FAST I will fan and medal what is -7x - 6 = less than 43 solve for X

OpenStudy (gorv):

-7x-6=<43 move 7x on right side -6=<43+7x move 43 on left -6-43=<7x -49=<7x -49/7=<x -7=<x